OpenStudy (anonymous):

A baseball is popped straight up into the air and has has a hang time of 6.25s. Determine the height to which the ball rises before it reaches it's peak

OpenStudy (zpupster):

half time going up and half time coming down so 6.25/2=3.125s from motion formula change in distance(d) = -1/2at^2 - vFinalT well vFinalt will = 0 going from + to - at the peak and a is gravity which is -9.8m/s^2 or 32.2ft/s^2 lets plug the values in d= -1/2at^2 d=(-1/2)(-9.8m/s^2)(3.125s)^2=?